# Post-detection inference for sequential changepoint localization

### v2

@@ B.1 Known pre-change and composite post-change
For weighted CUSUM or SR detectors (defined in (3.15),(3.16)), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{(\tau-j+1)\times(\theta-\theta_{0})\}(\theta/\theta_{0})^{\sum_{i=j}^{\tau}X_{i}}$ increases as $X_{i}$’s increases and hence both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
B.2 Composite pre- and post-change
@@
$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise. }\end{cases}$ (B.3)
Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (4.8), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.

### v3

@@ B.2 Another concrete example: Poisson rate change problem
For weighted CUSUM or SR detectors (defined in Remark 3.3.), finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For example, consider $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{(\tau-j+1)\times(\theta-\theta_{0})\}(\theta/\theta_{0})^{\sum_{i=j}^{\tau}X_{i}}$ increases as $X_{i}$’s increases and hence both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta\neq\theta_{0}\}$, if we employ Gamma weights (it is a conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution.
Appendix C Computationally and statistically efficient subroutines in Algorithm 3
Let $L_{t}^{n}(\theta_{0},\theta_{1})$ be the likelihood of data up to time $n\geq t$ under $H_{0,t}$. Then,
$L_{t}(\theta_{0},\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (C.1)
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})}.$ (C.2)
$M_{t}^{(n)}(X_{1},\cdots,X_{n}):=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{0,1:i-1},\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (C.3)
Note that the test statistic $M_{t}$ for testing the null hypothesis $H_{0,t}$ reduces to
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{0,1:\hat{T}-1},\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{0,1:t-1},\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau\text{ or }\tau=\infty.\\ \end{cases}$ (C.4)
As in Appendix B, we assume that we have a fixed element $\theta^{\prime\prime}$ in the parameter space and a function $G$, such that $X\sim F_{\theta^{\prime\prime}}$ implies $G(X,\theta)\sim F_{\theta},\forall\theta\in\Theta_{0}\cup\Theta_{1}$. Therefore, for each $j=1,\cdots,B$, we generate data sequence $X_{1}^{j},X_{2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{\theta^{\prime\prime}}$ only once. Then, for each $t\in\{1,\cdots,\tau\}$ and $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, we obtain sequences $\{Y_{n}^{t}(X_{n}^{j},\theta,\theta^{\prime})\}_{n}$ having joint distribution $\mathbb{P}_{\theta,t,\theta^{\prime}},$ by applying the function $G(.,\theta)$ to $X_{1}^{j},\cdots,X_{t-1}^{j}$ and $G(.,\theta^{\prime})$ to $X_{t}^{j},X_{t+1}^{j},\cdots$, i.e.,
$Y_{n}^{t}(x,\theta,\theta^{\prime}):=\begin{cases}G(x,\theta),&\text{if }n<t\\ G(x,\theta^{\prime}),&\text{ otherwise. }\end{cases}$ (C.5)
Note that, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta,\theta^{\prime})\}_{n}$ across all $(\theta\times\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
Let us now fix $t\in\{1,\cdots,\tau\}$. Computing the detection time $\tau^{j}_{\theta,\theta^{\prime}}$ for each $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$, and finding whether $\text{there exists }(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$ such that $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})$ are both computationally infeasible, unless $\mathcal{S}_{t-1}$ and $\mathcal{S}^{\prime}_{\tau-t+1}$ are a finite sets.
$U_{t,L}^{j}:=\begin{cases}-\infty,\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t\leq t_{2}^{j}\leq L\\ \infty,\leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L.\end{cases}$ (C.6)
And for $L=\infty$,
$U_{t,\infty}^{j}:=\begin{cases}-\infty,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty\\ \displaystyle\max_{\max\{t_{1}^{j},t\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{t^{\prime}}(Y_{1}^{t}(X_{1}^{j},\theta,\theta^{\prime}),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta,\theta^{\prime});t^{\prime}),\\ \leavevmode\nobreak\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$ (C.7)
We consider the following confidence set:
$\mathcal{C}^{\prime\prime}_{1}=\{t\in\{2,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},U_{t,L}^{1},\cdots,U_{t,L}^{B})\}\text{ and}$ (C.8)
Observe that $\mathcal{C}^{\prime\prime}_{1}\supseteq\mathcal{C}_{1}$, where $\mathcal{C}_{1}$ is the set produced by Algorithm 3, since
Corollary C.1.
$\mathcal{C}^{\prime\prime}_{1}$ defined in (C.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.5.
In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix C of the Supplementary Material.
C.1 A concrete example: Gaussian mean shift problem
$\mathcal{S}_{t-1}=\left[\bar{X}_{1:t-1}-\frac{q_{\gamma/2}}{\sqrt{t-1}},\bar{X}_{1:t-1}+\frac{q_{\gamma/2}}{\sqrt{{t-1}}}\right]\cap\Theta_{0},$ (C.9)
and the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.9).
$Y_{n}^{j}(\theta,\theta^{\prime})=\begin{cases}\theta+\epsilon_{n}^{j},\text{ if }n<t\\ \theta^{\prime}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (C.10)
We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in (C.4), with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Note that, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly even with $L=\infty$ for some specific detection algorithms; see Remark C.2. For $t_{2}^{j}<t,U_{t,\infty}^{j}=0$ and for $t_{2}^{j}\geq t,U_{t,\infty}^{j}$ can be simplified for our Gaussian setting as below:
$\begin{cases}\displaystyle\sum_{k=i}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}-\frac{t^{\prime}-t+1}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{t^{\prime}}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{i:t^{\prime}}^{j}+\frac{t-i}{t^{\prime}-i+1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i<t,\\[8.0pt] \text{constant},&\text{if }i=t,\\[8.0pt] \displaystyle\sum_{k=1}^{t-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}-\frac{i-t}{i-1}(\theta_{1}-\theta_{0})\right)^{2}\\ \quad+\sum_{k=t}^{i-1}\left(\epsilon_{k}^{j}-\bar{\epsilon}_{1:i-1}^{j}+\frac{t-1}{i-1}(\theta_{1}-\theta_{0})\right)^{2}+\text{constant},&\text{if }i>t,\end{cases}$ (C.11)
over $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Here, constant means that it does not involve $\theta,\theta^{\prime}$. Note that $\delta_{t^{\prime}}^{i}:=\theta_{1,t^{\prime}}^{i}-\theta_{0,t^{\prime}}^{i}$ have simple closed-form expressions since (C.1) is a quadratic function of $\theta_{1}-\theta_{0}$. Then, we can take $(\theta_{0,t^{\prime}}^{i},\theta_{1,t^{\prime}}^{i})$ to be any pair $(\theta,\delta_{t^{\prime}}^{i}+\theta)\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$.
Remark C.2.
[…]
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.12)
$\tau_{\text{wsr-ripr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (C.13)
where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For example, consider $\mathcal{P}_{0}=\{N(\theta,1):\theta\leq a\},\mathcal{P}_{1}=\{N(\theta,1):\theta\geq b\}$, for some $b\geq a$. Observe that for all $\theta\in\Theta_{1},f_{\theta}^{*}=f_{a}$, and $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}=\exp\{\frac{1}{2}(\theta-a)\times\sum_{i=j}^{n}(2X_{i}-a-\theta)\}$ increses as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore,
The Gaussian example we provided serves as a concrete and illustrative case to clarify the methodology. Note that $\mathcal{C}^{\prime\prime}$ in (C.8) is computable if similar conditions as listed in Section B.1 are met. For instance, one can check that these conditions hold for analogous Laplace and exponential scale distributions. Another concrete example with the Poisson distribution is in Section G.3.
C.2 Another concrete example: Poisson rate change problem
@@
$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise. }\end{cases}$ (C.14)
Using the binomial thinning property of Poisson distribution, it is straightforward to verify that $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta)$, for $n<t$, $Y_{n}^{j}(\theta,\theta^{\prime})\sim\text{Pois}(\theta^{\prime})$, for $n\geq t$ and $(\theta,\theta^{\prime})\in\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau};\tau)$, as defined in earlier, with $\hat{\theta}_{0,1:t-1}=\frac{1}{t-1}\sum_{k=1}^{t-1}X_{k}$ and $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. For Poisson distribution as well, $t_{1}^{j},t_{2}^{j}$ are easy to compute exactly for the same detection algorithms, similar to the Gaussian setting discussed in Remark 4.5.